OpenStudy (anonymous):

A point charge Q_1 is distance r from the center of a dipole consisting of charges (+-) q_2 separated by distance s. The charge is located in the plane that bisects the dipole. At this instant, assuming r >>s, what is the magnitude of the torque on the dipole? Express this factor in terms of q_2, Q_1, s, pi, and r. torque= l*F

OpenStudy (anonymous):

F = \[F= \frac{ Q_1*q_2*s }{ 4*\pi*r^3}\] l = sSin (theta)

OpenStudy (anonymous):

Torque on a dipole of moment p in an electric field E is given by their cross product\[\tau =p\times E= pE sin\theta\]Here in the case it is given that\[\theta =90^o\]So \[Sin\theta =1\]\[\tau =p.E\]and the dipole moment p is defined as \[p= s.q\]The electric field in question is of the charge Q at point 'r' away. Plug the expressions of this E and p (from above) into the formula for torque. It'll give the answer.